40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1694     Entry time: Mon Aug 1 18:47:58 2016
Author: awade 
Type: Summary 
Category: PLL 
Subject: Present electronic configuration of the PLL loop 

Post documenting the set out of the PLL electronics and cable routing.

We were having trouble with the PLL electronics in obtaining a beat note error signal and lock. We can see a beat note of 17 dB clearance above the DN level on the spectrum analyzer.   This post is to document the present configuration and the powers present at the mixer before I start to troubleshoot (aka pull it apart).

Schematic of the PLL electronics stage attached below.  Settings for the Marconi and SR560 are also below. I would noted that there is a low pass filter on the output of the phase detector (which I assume is a mixer) that extracts the (omega_LO - omega_sig) but has nowhere for the dumped (omega_LO + omega_sig) signals to go.  I'm guessing the non-DC signals are getting reflected directly back into the ZRPD-1+. I can't imagine that is great if back reflections start to form parasitic competing signals with the beat note signal.

PLL Loop electronics

Marconi 2023A
Carrier Frequency = 75.526 MHz (close to PD beatnote)
RF Level = +13 dBm
FM Devn 10.0 kHz (set by Ext DC)

SR560
Filter cutoff = DC
Coupling = DC
Source  = A (B port terminated with 50 ohm)
Inv = On
Gain Mode = Low noise
Gain  = 50
 

The spec sheet for the ZRPD-1+ indicates that it is a Level 7 mixer with a damage threshold at 50 mW, so the +13 dBm probably isn't a risk but may be saturating the mixer.  I don't have a calibrated power level for the RF PD input, but it is low (definitely not +7 dBm).  At this stage we either need to improve the beat note or implement some kind of amplification so that both the PD-RF input and LO inputs into the phase detector are matched at the ideal +7dBm level. 

At this stage both PDH control loops are ringing with DC signals that are very angular (and about 10% of the total DC signal). The angluarness would indicate that they are probably both railing somewhere or converting an control-error signal to amplitude modulations.  Previously we have diagnosed this back to imperfect polarization input into the fast control BB-EOMs.  I checked both North and South path and the inputs to the all EOMs (the 14 MHz and fast control) are very well aligned to 's': modulator polarization is not the issue. Antonio had this very well optimized form yesterday.

Tuning laser temperature around on the north path, it is apparent that there are some new/stronger HOMs around the region of our TEM00 modes. This is only a qualitative observation. It is possible that over time our mounts have drifted and the alignment is off.  This is now on my list of things check and to tune up.  

At this stage the PDH loop issues need to be resolved as they are impeding efforts to reliably form an observable beat note.

The process of unlocking and re-configuring for alignment and other types of diagnostic measurements is cumbersome.  There are BNCs popping out of the experiment everywhere that are also mostly unlabeled.  It would be much nicer to have all the useful control signals and PD signals coming out at a central point, preferably next to the computer interface and electronics rack where we lock and tune gains/temperatures.  I will gradually start migrating the cabling over to that side of the table and then install labeled 'patch lines' that are routed to mounted BNC panels around the table. This way there is a central location close to where we perform locking etc but it is relatively easy to patch through signals and oscilloscope trigger lines for temporary oscilloscopes situated around the table. This also means that consistent labeling can be applied.

It is not a great idea to label the cables with label-writer tape or color code them with electrical tape.  From experience the glues in these breakdown in 2-5 year timeframes either flaking off or leaving a sticky residue that is almost impossible to remove with solvents that don't also dissolve the BNC cable plastic. Heatshink can be good as it can be cut off and doesn't have glue. Otherwise a cable tied labels work well. We have some of these.

It also distresses me a little that BNCs cables are routing HV around the table without a clear system for distinguishing these cables. By Murphy's law these will one day be plugged into something sensitive. Some of them are labeled by are not marked clearly as HV. If I can find some red heat-shrink that fits over the ends of the BNC cables I might start labeling HV lines with a consistent markings to lower the risk of mix ups. 

 

 

Attachment 1: 20160801_ElectronicLayoutOfPLLLoop.pdf  78 kB  | Hide | Hide all
20160801_ElectronicLayoutOfPLLLoop.pdf
ELOG V3.1.3-